a relief aeroplane is flying at a constant height of 1960 M with the speed 600 km per hour above the ground towards a point directly over a person struggling in a flood water at what angle of sight with the vertical should the pirate release a survival kit to reach the person in the flood water

Grade:10

1 Answers

Arun
5 years ago
Dear Keshav
 
Let @ is the angle of sight
 
then
 
tan@ = x / h
 
horizontal speed of kit = vx = 600 km/hr = 166.66 m/sec
 
vertical speed v 
v^2 – u^2 = 2as
 
v^2 = 2 * 10 * 1960 
 
v = 197.98 m/sec
 
tan@ = 197.98 / 166.66 = 1.18
 
@ = tan^-1 (1.18)